OpenStudy (anonymous):

Find the rate of change of the area of a circle per second with respect to its radius r when r = 5 cm.

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Differentiate \(\pi r^2\) with respect to \(r\).

Parth (parthkohli):

When you find \(f'(r)\), find \(f'(5)\)

OpenStudy (anonymous):

Thanks

Parth (parthkohli):

The answer turns out to be the circumference.

OpenStudy (badhi):

Actually shouldn't it be like this, $$A=\pi r^2$$ Since rate of change is associated with 'time' differentiation with respect to 't'- time, $$\frac{dA}{dt}=\pi\frac{dr^2}{dt}=2\pi r\frac{dr}{dt}$$ rate of change in radius should be given